Please answer with steps and laws/rules used. Thank you!1. Find an equation of the tangent line to the graph of y = (5x at (0,1).i/pi. 2. Find an equation of the tangent line to the
Answer #1

2 y = 5x 7 dly, 505X (Chain Rule) 00) 52-5 og of tengent y = 5(2-0) = + 5x | - 29 22 = 25 differentiate, 24*24 l at (2,4), dy